# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v2

@@ 3.2 SPike-Aware Consistency Enhancement
The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data. When the feature maps of augmented samples are highly similar, the model becomes less sensitive to variations at test time, thereby improving its ability to generalize to new, unseen, and unlabeled data. To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$. Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ denotes the local feature vectors associated with channel $c$ of the augmented sample $\mathbf{x}_{i}$. Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$. This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively. To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
The similarity measures the degree of overlap between the feature distributions, with higher values indicating greater consistency in spike-based representations across augmentations. Based on this similarity measure, we define the loss function for model adaptation given as
$\mathcal{L}(\theta;\mathbf{x})\triangleq\sum_{1\leq j<i\leq M}\left(1-\bar{\mathcal{S}}(i,j|\mathbf{x})\right).$ (5)
Then, we adapt the model using stochastic gradient descent (SGD) for a single iteration with the learning rate $\eta$. By minimizing this loss function, we enforce consistency across the feature maps of different augmentations. This reduces the variability induced by augmentations, ensuring that the model learns stable and robust feature representations. Furthermore, by promoting feature consistency, this approach serves as a regularization mechanism, preventing over-fitting to specific augmentations and encouraging the model to focus on more fundamental and generalizable features.
While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure. Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) in Appendix). However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
Finally, we predict the classification via an updated model.
Discussion
In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally. Since channels may have different active neurons, directly comparing the entire feature map fails to account for this variability. Instead, computing the similarity per channel and averaging the results focuses on the stability of feature representations within each channel. This isolates the contributions of active neurons, ensuring the similarity measure reflects meaningful changes rather than noise. It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels. Our experiments in subsection 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach. Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.
@@ 4 Experimental Results
The experiments were conducted on six benchmark datasets for out-of-distribution generalization: CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], and ImageNet-A [18]. The first three datasets include 15 corruption types across four categories and five severity levels, while the three ImageNet-based datasets capture real-world domain shifts, such as distributional discrepancies and semantic variations. Additionally, we evaluated on the neuromorphic dataset DVS Gesture-C [20], which introduces six corruption types. Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[24], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[27]; 3) Spike-driven Transformer V3 [44], containing 19M parameters; and 4) SNN-ConvLSTM. All models were pre-trained on their respective source domains to ensure fair evaluation on datasets with domain shifts.
@@
We compared our approach with models pre-trained on clean datasets and evaluated without any test-time adaptation (referred to as No Adapt). Additionally, since there are few TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [47] and SITA [23], as baselines. These methods represent the current state-of-the-art for source-free and single-instance TTA, respectively. MEMO enforces output consistency among augmented samples, making it naturally compatible with SNNs by directly adjusting their final outputs. SITA, designed for models with BN layers, was adapted to SNN-VGG with BNTT by updating BN layer parameters during TTA.
@@
To maximize the effectiveness of our proposed method, all experiments conducted on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C , and DVS Gesture-C were performed with the highest severity level (level=5). In line with our setting and practical scenarios, the test batch size for all experiments was fixed to 1, and the model was updated only once per test sample. Following MEMO’s experimental setup, we applied AugMix [16] augmentation to generate diverse augmented samples, using a batch size of 32 for CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C, and 64 for the ImageNet series. Please refer to Appendix for more details.
@@ 4.1 Main Results
In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO and SITA. To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using SNN-VGG, on CIFAR-10-C with SNN-ResNet11, on ImageNet series with Spike-driven Transformer V3, and on DVS Gesture-C with SNN-ConvLSTM, to demonstrate the broad applicability of our proposed method.
@@
The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Table 1, highlight the effectiveness of SPACE. Key observations include the following: 1) Superior Performance Across Methods: SPACE consistently outperforms No Adapt, SITA, and MEMO in average accuracy across all datasets under both VGG and ResNet architectures. 2) Broad Applicability: SPACE demonstrates effectiveness across different CNN structures, showcasing its generalizability. 3) Robustness on Complex Datasets: A particularly notable result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type. This highlights its robustness in handling complex datasets like Tiny-ImageNet-C. 4) Limitations of SITA on SNNs: The SNN-ResNet does not include BN layers, rendering SITA inapplicable. Furthermore, SITA performs worse than No Adapt, indicating its ineffectiveness for SNNs. This may be attributed to the spike-based transmission in SNNs: directly adjusting BN parameters without gradient guidance can disrupt spike information, leading to significant performance degradation. 5) Comparison with MEMO: While SPACE generally outperforms MEMO and SITA, MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs. 82.88% for SPACE) and CIFAR-100-C (59.29% for MEMO vs. 58.96% for SPACE). These minor differences suggest MEMO may have an advantage under certain corruption types, but they do not detract from SPACE’s overall superiority.
@@
To explore the real-world impact of SPACE, we conducted tests on diverse datasets, including ImageNet-V2, ImageNet-R, and ImageNet-A. Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-Event Transformer V3. As shown in Figure 3, SPACE consistently outperforms MEMO, achieving a 1.80% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
@@
SNNs are inherently well-suited for processing event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency. The DVS-Gesture dataset [1] provides an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures. To further explore potential network architectures, we incorporated ConvLSTM. As shown in Table 2, our approach achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
@@ 4.2 Ablation Studies
(a) Local feature maps based SPACE outperforms global feature maps based SPACE.
(b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
Figure 4: Results of ablation study.
@@
We compared the performance of local versus global feature map similarity for our proposed method. The results in 4(a) clearly illustrate the advantage of using similarity of local feature maps over global feature maps. While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency. As shown in 4(a), the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets. This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation. These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
@@
To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset. As shown in 4(b), we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain. The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process. Specifically, the "After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples. This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
5 Conclusion and Limitation
@@
In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs. By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations. Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures, demonstrating robustness to domain shifts. However, SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios. Future work will explore alternative optimization strategies to expand its practicality for neuromorphic hardware and energy-efficient AI.
Appendix A Kernel-Based Consistency Regularization

### v3

@@ 3.2 SPike-Aware Consistency Enhancement
Our goal is to make the model focus on invariant, task-relevant features rather than noise or minor transformations introduced by augmentations or distribution shift. If the feature maps of augmented views are highly similar, the model becomes less sensitive at test time and generalizes better to unseen, unlabeled samples. Concretely, given the spike-aware feature map $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ of an augmented view $\mathbf{x}_{i}$, we compute channel-wise local vectors $\mathbf{F}_{c}(\mathbf{x}_{i})\in\mathbb{R}^{D}$ and normalize them across spatial locations with a softmax to obtain a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})\in\Delta^{D-1}$. This normalization emphasizes salient regions and suppresses noise, yielding a stable target distribution. We then measure similarity between two augmented views $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ by the average channel-wise inner product
$\bar{\mathcal{S}}(i,j\mid\mathbf{x})\;=\;\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\,\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
which lies in $[0,1]$ and attains 1 if and only if the per-channel spatial distributions match exactly. Based on this, we adapt the extractor parameters by minimizing the loss function
$\mathcal{L}(\theta_{E};\mathbf{x})\;\triangleq\;\sum_{1\leq j<i\leq M}\bigl(1-\bar{\mathcal{S}}(i,j\mid\mathbf{x})\bigr),$ (5)
using a single SGD step with learning rate $\eta$. Minimizing $\mathcal{L}$ enforces consistency across augmentations, regularizing the model to learn stable, robust representations and preventing overfitting to any specific view. Finally, predictions are obtained with the adapted model.
Theoretical Motivation
Our objective instantiates consistency regularization at inference time: under semantics-preserving, bounded-severity augmentations (e.g., AugMix [19]), we enforce invariance in internal representations across transformations. Information-theoretically, aligning $\mathbf{F}$ (and thus $\mathbf{P}$) suppresses augmentation-specific variability while preserving identity-relevant content, approximating an information bottleneck that favors invariant features; geometrically, when perturbations push a sample off the learned manifold, adaptation pulls it toward the local manifold spanned by its augmented views, effectively denoising the representation. In SNNs, shifts often manifest as event-rate changes and timing jitter that nudge membrane potentials $U$ around the threshold $U_{th}$, making spike generation highly sensitive; by aligning time-resolved features across augmentations, we reduce the variance of $U$ trajectories and spike counts, enlarge the effective margin $|U-U_{th}|$, and stabilize and spike timing dynamics. Consequently, robustness arises from coupling semantic consistency with temporal stabilization within the SNN-TTA framework.
We also explored richer feature relations via kernel embeddings and Maximum Mean Discrepancy (MMD) (Appendix B). In practice, the above linear-kernel objective already captures the essential distributional overlap with lower variance and cost, and stronger kernels did not bring noticeable gains while adding overhead. More discussion is presented in Appendix A.
@@ 4 Experimental Results
Table 1: Performance of different model architecture under domain shift. Values are Top-1 accuracy (%); Acc. Loss = Clean - Shifted. “Shifted” denotes evaluation on shifted variants of each dataset; exact configuration described in Appendix C. SNN and ANN models share backbone and scale and are trained with the same protocol.
Architecture Dataset ANN SNN
Clean Shifted Acc. Loss Clean Shifted Acc. Loss
VGG9 CIFAR-10 90.53 77.25 13.28 90.61 67.86 22.74
VGG11 CIFAR-100 70.86 48.32 22.54 69.50 41.35 28.14
VGG11 Tiny-ImageNet 61.24 19.76 41.48 58.36 14.56 43.80
Transformer ImageNet 85.16 75.49 9.67 79.60 67.64 11.96
@@
The experiments were conducted on six benchmark datasets for out-of-distribution generalization: CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [18], ImageNet-V2 [42], ImageNet-R [20], and ImageNet-A [21]. The first three datasets include 15 corruption types across four categories and five severity levels, while the three ImageNet-based datasets capture real-world domain shifts, such as distributional discrepancies and semantic variations. Additionally, we evaluated on the neuromorphic dataset DVS Gesture-C [23], which introduces six corruption types. Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[29], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[32]; 3) Spike-driven Transformer V3 (19M parameters) with 4 inference steps [52]; and 4) SNN-ConvLSTM with 32 inference steps. All models were pre-trained on their respective source domains to ensure fair evaluation on datasets with domain shifts.
@@
We compared our approach against models pre-trained on clean datasets and evaluated without any test-time adaptation (henceforth referred to as No Adapt), which serves as a lower-bound reference for robustness under distribution shift. Given the limited availability of TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [56] and SITA [28], as baselines. These methods represent the current state-of-the-art for source-free and single-instance TTA, respectively. MEMO enforces output consistency across augmented samples, making it inherently compatible with SNNs by directly adjusting their final outputs. SITA, originally designed for models with BN layers, was adapted for SNN-VGG with BNTT by updating the BN layer parameters during TTA. Additionally, we included RoTTA [54] and the recent DeYO [34] as baselines, despite their reliance on batch data. To ensure a fair comparison, we applied the same augmentation strategy as SPACE to construct a pseudo-batch for each sample.
@@
To maximize the effectiveness of our proposed method, all experiments conducted on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C , and DVS Gesture-C were performed with the highest severity level (level=5). In line with our setting and practical scenarios, the test batch size for all experiments was fixed to 1, and the model was updated only once per test sample. Following MEMO’s experimental setup, we applied AugMix [19] augmentation to generate diverse augmented samples, using a batch size of 32 for CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C, and 64 for the ImageNet series. Please refer to Appendix C for more details.
4.1 Evaluation of Robustness of SNNs under Domain Shift.
To further substantiate the observation that SNNs are particularly vulnerable to domain shifts, we conducted a controlled comparison between SNNs and ANNs that share the same backbone family and model scale, trained under identical protocols, and evaluated on shifted variants of standard benchmarks (details in Appendix C). As summarized in Table 1, SNNs consistently incur larger accuracy drops than their ANN counterparts across CIFAR-10, CIFAR-100, Tiny-ImageNet, and ImageNet. These empirical results motivate the development of robust, SNN-specific TTA strategies tailored to spiking computation.
@@ 4.2 Main Results
In this section, we compare SPACE’s performance with four existing TTA methods. To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using SNN-VGG, on CIFAR-10-C with SNN-ResNet11, on ImageNet series with Spike-driven Transformer V3, and on DVS Gesture-C with SNN-ConvLSTM, to demonstrate the broad applicability of our proposed method.
@@
The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Table 2, highlight the effectiveness of SPACE. Key observations include: 1) Optimal Performance: SPACE consistently outperforms all baselines in average accuracy across all datasets under both VGG and ResNet architectures. 2) Broad Applicability: SPACE demonstrates effectiveness across different CNN structures, showcasing its generalizability. 3) Robustness on Complex Datasets: A particularly notable result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type. This highlights its robustness in handling complex datasets like Tiny-ImageNet-C. 4) Vs. RoTTA: RoTTA’s memory bank cannot be effectively utilized in this single-sample setting, leading to limited performance gains. 5) Vs. SITA: The SNN-ResNet does not include BN layers, rendering SITA inapplicable. Furthermore, SITA performs worse than No Adapt, indicating its ineffectiveness for SNNs. This may be attributed to the spike-based transmission in SNNs: directly adjusting BN parameters without gradient guidance can disrupt spike information, leading to significant performance degradation. 6) Vs. DeYO: DeYO relies on smooth probability outputs for confidence comparison, which is incompatible with SNN’s discrete, temporal spike-based outputs, introducing significant errors. Also, DeYO only works normalization layer. 7) Vs. MEMO: MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs. 82.88% for SPACE) and CIFAR-100-C (59.29% for MEMO vs. 58.96% for SPACE). The minor edge cases hint at MEMO’s advantage under certain corruption types, yet they don’t diminish SPACE’s overall superiority.
@@
To explore the real-world impact of SPACE, we conducted tests on diverse datasets, including ImageNet-V2, ImageNet-R, and ImageNet-A. Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-driven Transformer V3. As shown in Figure 3, SPACE consistently outperforms all baselines, achieving at least 1.59% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
@@
SNNs are well-suited to event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency. The DVS-Gesture dataset [1] is an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures. To further explore potential network architectures, we incorporated ConvLSTM. BN-dependent SITA and DeYO is not suitable for this architecture. As shown in Table 3, SPACE achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
@@ 4.3 Ablation Studies
Figure 4: a) Local feature maps based SPACE outperforms global feature maps based SPACE. b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
@@
Definitions. Let “global feature maps" refers to treating the entire feature map of size $C\times H\times W$, as a single entity and measuring similarity on this full representation. Let the “local feature map" considers each spatial position in the $H\times W$ dimension independently, which is adopted to calculate the similarity and averaging across the $C$ channels in the main paper.
We compared the performance of local versus global feature map similarity for our proposed method. The results in Figure 4a clearly illustrate the decisive advantage of using similarity of local feature maps over global feature maps. While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency. As shown in Figure 4a, the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets. This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation. These findings underscore the distinct roles of each channel in SNNs and further highlight the practical limitations of comparing entire feature maps.
@@
further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset. As shown in Figure 4b, we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain. The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process. Specifically, the “After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples. This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
Table 4: Comparison of computational cost on CIFAR-10-C with SNN-VGG9 model.
Method RoTTA DeYO SITA MEMO SPACE
GFLOPs 160.67 318.07 161.48 161.48 158.11
Accuracy 67.79% 67.99% 66.41% 69.20% 71.03%
Analyses of Computational Cost
To evaluate computational efficiency, we compared the GFLOPs of different baselines with the same augmented batch size, and SPACE achieved the lowest GFLOPs, demonstrating its efficiency (refers to Table 4). Regarding the additional cost of optimizing spike-behavior consistency, this step is performed only once per input in single-sample setting, without requiring whole model backpropagation or iterative updates.
Effect of Augmentation Quantity
In the main paper, we follow MEMO [56] to determine the number of augmentations $M$. To assess $M$’s effect, we evaluated the average accuracy and the evaluation time per test point under the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$. As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement. For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly. Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
No Adapt Number of Augmentation
2 4 8 16 32 64 128
Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
5 Conclusion
@@
In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs. By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, directly addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations. Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series distribution shifted datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures and backbones, demonstrating robust, practical resilience to real-world domain shifts.
Acknowledgments and Disclosure of Funding
This work was supported by Hong Kong Innovation and Technology Fund (ITF) MSRP grant ITS/018/22MS.
Appendix A Discussion
Selection of Alignment Layer
We compared aligning the classifier’s feature map with aligning extractor layers used in the main paper. Empirically, classifier-level alignment occasionally degrades performance on several corruptions, whereas extractor-level alignment is consistently more stable. We attribute this to: 1) classifier features being highly task-specific and high-curvature—small per-sample TTA steps can move decision boundaries and induce catastrophic forgetting [5]; 2) lower signal-to-noise at the classifier under pseudo-label or consistency noise, which amplifies domain-specific biases; and 3) greater adaptation overhead when updating deeper, larger parameter blocks. In contrast, channel-wise alignment at mid-level layers promotes augmentation-invariant structure while keeping the classifier intact, offering a better stability–plasticity trade-off with lower compute.
Potential Information Loss in Equation 4
We aggregate spikes over time, convert discrete spike counts into continuous probability values via a softmax function, then calculate the similarity. This transformation is the intentional discarding of the absolute magnitude of spike counts to focus on the relative importance of neurons within a channel. By applying softmax, we normalize away noisy fluctuations in absolute firing rates and create a stable spatial saliency map. Aligning these maps forces the model to learn which spatial features are consistently most important, directly encouraging the learning of invariant representations. In addition, while our alignment objective operates on this spatially-focused abstraction, it simultaneously preserves and refines the inherent temporal characteristics of the SNN. The underlying model remains a fully temporal processor, and the spike counts are the direct result of its complex dynamics. By demanding consistency in the outcome (the count-based saliency map), our method implicitly forces the network to stabilize the cause (the spike generation process). This creates a powerful regularization pressure that reduces temporal jitter and enhances the robustness of spike timing, all without the computational cost and overfitting risks associated with direct, spike-for-spike temporal alignment.
Why not cosine or KL similarity
We compute similarity as a per-channel dot product over softmax-normalized spike counts, which is bounded, symmetric, and well-defined even when raw counts are zero. Cosine similarity requires $\ell_{2}$ normalization and is numerically fragile under SNN sparsity: many channels yield zero or near-zero vectors, causing division-by-zero or unstable gradients; adding an $\varepsilon$ fixes numerics but introduces a sensitive hyperparameter and slightly degrades performance in our trials. KL divergence, $\mathrm{KL}(\mathbf{P}\|\mathbf{Q})$, is asymmetric and becomes ill-conditioned when entries approach zero (common with peaked softmax under corruptions), leading to exploding losses unless additional smoothing is used—again adding tuning burden. A symmetrized variant (e.g., JSD) alleviates this but increased compute without improving over the simple dot product. Given our per-sample TTA regime, the dot product on per-channel probability maps offers the best robustness–efficiency trade-off, avoiding normalization pitfalls while retaining discriminative alignment.
Feasibility of Deployment
On neuromorphic platforms like SpiNNaker2 [14], the data pre-processing unit and neural network inference module are typically separate. Batch data augmentation can be performed in the pre-processing stage, outside the neuromorphic core, before feeding data into the system for inference. This design allows augmentation to be handled efficiently without impacting the neuromorphic computing core, making our method feasible on such platforms.
Limitation
Proposed SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios. Future work will explore alternative optimization strategies to expand its practicality for more neuromorphic hardware and energy-efficient AI.
Appendix B Kernel-Based Consistency Regularization
Method Noise Blur Weather Digital Average